OpenStudy (anonymous):

the formula S=16t^2 is used to approximate the distance S, in feet, that an object falls freely from the rest in t seconds. The height of a building is 1278 feet. How long would it take for an object to fall from the drop?

OpenStudy (mertsj):

Replace s with 1278 and solve for t

OpenStudy (anonymous):

i got -3/4 square root of 143 and postive 3/4 square root of 143 ??? but the answering im looking for is suposed to be a decoimal

OpenStudy (anonymous):

decimal

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i got -3/4 square root of 143 and postive 3/4 square root of 143 ??? but the answering im looking for is suposed to be a decoimal

OpenStudy (mertsj):

\[1278=16t^2\] \[79.875=t^2\] t=8.9

OpenStudy (anonymous):

no thats not right

OpenStudy (mertsj):

Then what is the answer?

OpenStudy (anonymous):

i dont know but i triedthat and it didnt work.

OpenStudy (anonymous):

did you mean the answer is 8.9 times 8.9

OpenStudy (mertsj):

Maye you need more decimal places. How about 8.94 or 8.937

OpenStudy (mertsj):

No. The answer is 8.9

OpenStudy (anonymous):

wow it was 9.0 that was my fault. can u help with a similar one? S=16t^2 1479 feet

OpenStudy (anonymous):

wow it was 9.0 that was my fault. can u help with a similar one? S=16t^2 1479 feet

OpenStudy (mertsj):

same thing. 1479=16t^2 Divide both sides by 16 and take the square root of that answer.

OpenStudy (anonymous):

i got 9.6

OpenStudy (mertsj):

Me too.

OpenStudy (anonymous):

Can you help me with another question?

OpenStudy (mertsj):

post it

OpenStudy (anonymous):

a 15 ft long ladder leans on a wall, the bottom of the ladder is 9 ft from the wall.if the bottom is pulled out 4 ft farther from the wall, how far does the top of the ladder move down the wall

OpenStudy (anonymous):

round to the nearest thousanth as needed

OpenStudy (mertsj):



OpenStudy (mertsj):

Use the Pythagorean theorem to find x

OpenStudy (mertsj):